### v2

# Recommendation aided Caching using Combinatorial Multi-armed Bandits

Abstract
We study content caching with recommendations in a wireless network where the users are connected through a base station equipped with a finite-capacity cache. We assume a fixed set of contents with unknown user preferences and content popularities. We can recommend a subset of the contents to the users which encourages the users to request these contents. Recommendation can thus be used to increase cache hits. We formulate the cache hit optimization problem as a combinatorial multi-armed bandit (CMAB). We propose a UCB-based algorithm to decide which contents to cache and recommend. We provide an upper bound on the regret of our algorithm. We numerically demonstrate the performance of our algorithm and compare it to state-of-the-art algorithms.
Index Terms:
caching, recommendation, CMAB, cache hit, UCB
I Introduction
As various intelligent devices, e.g., smart vehicles, home appliances, and mobile devices continue to evolve alongside a plethora of innovative applications, mobile wireless communications are witnessing an unprecedented surge in traffic. This surge is characterized by a significant amount of redundant and repeated information, which limits the capacity of both the fronthaul and backhaul links . Utilizing content caching at the network edge, including base stations, presents a promising strategy for addressing the rapidly growing demand for data traffic and enhancing user satisfaction.
Recently, recommendation mechanisms have demonstrated their effectiveness in enhancing the caching performance of conventional networks. By reshaping users’ content request behaviors, these mechanisms encourage edge caches to become more user-friendly and network-friendly [1]. Notably, studies by [18] and [8]revealed that 50% of YouTube video views are driven by recommendation systems, a figure that rises to 80% for Netflix [7]. However, independently deploying recommendation and caching strategies may not fully leverage their potential. This is because recommendations influence users’ content request behaviours, subsequently impacting caching decisions.
In our problem, we model the caching with a recommendation as a Combinatorial Multi-armed bandit (CMAB). In CMAB, the agent selects the subset of arms in each round and recieves the sum of rewards. The goal is to maximize this reward over a period of time. In our problem, each content corresponds to an arm in CMAB and we have to decide what contents to cache and what contents to recommend each time so that we can maximize the cache hit. In our problem users preference distribution i.e. users request distribution without recommendation is unknown. We can not learn this by observing the requests because recommendations influence users’ content request behaviours, subsequently impacting caching decisions. Caching decision also influences recommendation. Hence we need to solve recommendation and caching jointly.
I-A Related Works
In recent times, numerous research studies have suggested leveraging content recommendations to enhance caching efficiency. In references [2] and [1], recommendations are used to guide user requests towards cached content that aligns with their interests. Furthermore, recent developments have explored employing content recommendation to fulfill content requests by offering alternative or related content, as discussed in references [11], [12], and [3]. This method finds application in various domains such as video and image retrieval and entertainment-related services, as indicated in reference [11]. Considering the advantages of having recommendations in caching , there have been a lot of works on joint content caching and recommendation [13], [15], [1] and [2]. In [1], it was initially demonstrated that user preferences could be reshaped to enhance caching performance. Subsequent studies have further confirmed these findings, showing notable improvements in performance compared to traditional caching methods that do not integrate recommendation systems
Organisation of the paper
The rest of the paper is organised as follows. In Section (II) we present the system model for content caching with recommendation problem and formulate it as CMAB. In Section (III), we describe the algorithm to solve the problem and also provide upper bound on the regret. In Section (IV), we show the cache hit performance of our algorithm and also compare it with few existing algorithm. Section (V) concludes the section with some insight on future directions.
II System Model and Caching Problem
We consider a wireless network where the users are connected to a single base station (BS) which in turn is connected to content servers via the core network. The content providers have a set of $N$ contents, ${\cal N}=\{1,2,\cdots,N\}$ of equal size at the servers. There are $U$ users; ${\cal U}=\{1,2,3,...,U\}$ denotes the set of users. The BS has a cache where it can store up to $C$ contents. We assume a slotted system. The caching and recommending decisions are taken at the slot boundaries. We denote the caching vector by $\mathbf{Y}=\{y_{1},y_{2},y_{3},....,y_{N}\}$ where $y_{i}\in\{0,1\}.$ $y_{i}=0$ implies that the $i^{th}$ content is not cached and $y_{i}=1$ implies that it is cached. Similarly, we denote the recommending vector for user $u$ by $\mathbf{X_{u}}=\{x_{u1},x_{u2},...,x_{uN}\}$ where $x_{ui}\in\{0,1\}$. If $x_{ui}=1$ then the $i^{th}$ content is recommended to user $u$, else it is not recommended to user $u$. We assume that at most $R$ contents can be recommended to any user, where $R\leq C$. The reason for choosing $R\leq C$ is that users may ignore the recommendation if we recommend too many items. Hence, we recommend a smaller number of contents than those cached.
We aim to employ a recommendation strategy to enhance cache hit performance. A cache hit happens whenever requested content is stored in the cache. We assume that the base station (BS) does not know the users’ preferences, and the BS controls the recommendation system. The relationship between the recommendations and the user preferences is modelled as follows. We denote $p_{u}^{\text{pref }}(i)$ as the probability that user $u$ would request file $i$ in the absence of recommendations. We denote the $\mathcal{R}_{u}$ be the set of files recommended to user $u$. We assume that the recommendations to user $u$ induce a preference distribution $p_{u}^{\text{rec}}(\cdot)$; $0\leq p_{u}^{\text{rec }}(i)\leq 1,\forall i$, $p_{u}^{\text{rec }}(i)=0,\forall i\notin\mathcal{R}_{u}$, and $\sum_{i=1}^{R}p_{u}^{\text{rec}}(i)=1$. By jointly considering $p_{u}^{\text{pref }}(\cdot)$ and $p_{u}^{\text{rec }}(\cdot)$, the request probability of user $u$ for file $i$ is then modeled as
$p_{u}^{\text{req }}(i)=w_{u}^{\text{rec }}p_{u}^{\text{rec }}(i)+\left(1-w_{u}^{\text{rec }}\right)p_{u}^{\text{pref }}(i).$ (1)
Here $0\leq w_{u}^{\mathrm{rec}}\leq 1$ is the probability that user $u$ accepts the recommendations from the recommendation system. From (1), we can see that the final request distribution of the user is a convex combination of its original preference distribution and the preference distribution induced by the recommendations. Thus, the recommendation system shapes the probability distribution for file-requesting. We can see from (1) that the request probabilities for the recommended contents are boosted and the request probabilities for non-recommended contents are decreased. There are other ways of combining $p_{u}^{\text{rec}}(i)$ and $p_{u}^{\text{pref}}(i)$, e.g., $p_{u}^{\text{req }}(i)=\max\{p_{u}^{\text{rec }}(i),p_{u}^{\text{pref }}(i)\}$ [14], but a convex combination is the most general way of combining and has been used in most of the related works [2, 13, 16, 6, 4].
In our problem, we assume that the users select contents from the recommended list uniformly, i.e.,
$p_{u}^{\text{rec }}(i)=\begin{cases}1/R&\text{if }x_{ui}=1,\\ 0&\text{otherwise.}\\ \end{cases}$ (2)
This uniform distribution of recommended content has also been considered in many literature [13, 1]. We can also consider the case where the position of contents in the recommended list affects the probability of getting selected by the user, i.e. usually, the contents at the top have a higher probability of getting selected than those at the lower positions [10, 18, 8, 17]. In Remark 2, we justify that our analysis follows for the case when position in the recommended list impacts the request probability including (2).
Also, $w_{u}^{\mathrm{rec}}$ captures the percentage of time user $u$ accepts the recommendation. It has been reported that $50\%$ and $80\%$ of the requests come from recommendations on YouTube and Netflix, respectively [18, 8, 7]. But, it can, of course, differ among the users. Also, $w_{u}^{\mathrm{rec}}$ can change over time depending on the user’s interest. It is possible to estimate $w_{u}^{\mathrm{rec}}$, but we assume that $w_{u}^{\mathrm{rec}}$ is known. We would like to consider when $w_{u}^{\mathrm{rec}}$ is unknown in our future work.
II-A Content Caching Problem
In this section, we formulate the problem as a CMAB problem to maximize the cache hit. The sum of the cache hit probabilities across all the users is given by
$\sum_{u}\sum_{i}y_{i}p_{u}^{\text{req }}(i)$
$=\sum_{u}\sum_{i}y_{i}\bigg{[}w_{u}^{\text{rec}}\frac{x_{ui}}{R}+(1-w_{u}^{\text{rec}})p_{u}^{\text{pref}}(i)\bigg{]},$ (3)
where (3) follows from the request distributions (1) and Assumption (2). So our goal is to solve the following optimization problem to get the optimal caching and recommendations
$\max_{\mathbf{X},\mathbf{Y}}\sum_{u}\sum_{i}y_{i}$ $\bigg{[}w_{u}^{\text{rec}}\frac{x_{ui}}{R}+(1-w_{u}^{\text{rec}})p_{u}^{\text{pref}}(i)\bigg{]}$ (4)
subject to $\sum_{i}y_{i}\leq C$ (5)
$\sum_{i}x_{ui}\leq R\quad\forall u\in U$ (6)
$x_{ui}\in\{0,1\},y_{i}\in\{0,1\}\quad\forall u,i$ (7)
Constraint (5) accoutns for the cache capacity, and (6) for the number of recommended contents to a user. Constraint (7) indicates the contents cannot be partially cached or recommended
In Problem (4), user preference distributions in the absence of recommendation, $p_{u}^{\text{pref}}(i)$ for all $u,i$, are unknown. One way of solving this problem is to first estimate $p_{u}^{\text{pref}}(i)$ for all $u,i$, and then solve the optimization problem. But recommendations affect the number of requests, so we cannot directly estimate $p_{u}^{\text{pref}}(i)$. We can only estimate the request distributions $p_{u}^{\text{req}}(i)$. Also, we can only observe the requests of the cached contents and hence it is modeled as CMAB(i.e., in CMAB , only the rewards of the pulled arms are observed). This is different from full observation setting where we can observe the requests of all the contents irrespective of whether the content is cached or not. So in the next section, we introduce a UCB-based algorithm that estimates $p_{u}^{\text{req}}(i)$ for all $u,i$, and solves Problem (4).
III Algorithm Design
This section introduces an UCB-based algorithm, Algorithm 1, to solve Problem (4). We smartly choose the confidence interval incorporating $w_{u}^{rec}$. On a careful look at Problem  (4) we see that, for given $y_{i}$s, the first term is maximized when $x_{ui}=1$ only for those $i$ for which $y_{i}=1$. Hence, we should recommend cached contents only. Moreover, under this choice of $x_{ui}$s, the first term is independent of the choice of $y_{i}$s. To maximize the second term, we choose $Y^{*}$ as follows
$\quad\mathbf{Y}^{*}=\underset{\mathbf{Y}}{\operatorname{argmax}}\sum_{i}y_{i}\sum_{u}\left(1-w_{u}^{\text{rec }}\right)p_{u}^{\text{pref }}(i).$ (8)
Unfortunately, $p_{u}^{\text{pref }}(i)$ are unknown and we cannot estimate these as we explained in the previous section. If we knew $p_{u}^{\text{pref }}(i)$, the optimal solution would be to choose top $C$ contents with the highest $\sum_{u}\left(1-w_{u}^{\text{rec }}\right)p_{u}^{\text{pref }}(i)$. Let us denote the optimal set of contents by $\mathcal{Y}=\left\{i:y_{i}^{*}=1\right\}$ and the set of contents cached at time $t$ according to Algorithm 1 by $y(t)=\left\{i:y_{i}^{t}=1\right\}$. Now we define the regret for not caching the optimal contents $\mathcal{Y}$ as follows
$\operatorname{R}(T)$ $=T\sum_{u}\sum_{i}y_{i}^{*}p_{u}^{\text{req }}(i)-\sum_{t=1}^{T}\sum_{u}\sum_{i}y_{i}^{t}p_{u}^{\text{req }}(i)$
$=T\sum_{i\in\mathcal{Y}}\sum_{u}p_{u}^{\text{req }}(i)-\sum_{t=1}^{T}\sum_{u}\sum_{i\in y(t)}p_{u}^{\text{req }}(i)$
=$\stackrel{{\scriptstyle(a)}}{{=}}T\sum_{i\in\mathcal{Y}}p_{i}-\sum_{t=1}^{T}\sum_{i\in y(t)}p_{i}$ (9)
where in (a), we have used $p_{i}=\sum_{u}p_{u}^{\text{req }}(i)$. If we knew $p_{i}$s for all $i$, the optimal solution $\mathcal{Y}$ could also be given by following: arrange the contents in the decreasing order of $p_{i}$s and choose the top $C$ contents. In the absence of knowledge of $p_{i}$’s, we estimate these for all $i$ at each time. We can estimate $p_{i}$’s as follows. We use $Z_{i}(t)$ to denote the number of request for cached content $i$ at time $t$. We also define
$p_{i}(t)=\frac{{\sum_{s=1}^{t}Z_{i}(s)}}{n_{i}(t)}$ (10)
where $n_{i}(t)$ is the number of times content $i$ is cached.
Our UCB-based algorithm works as follows. We will estimate $p_{i}$’s $\forall i$. Then, construct the UCB index $U_{i}(t)=p_{i}(t)+D_{i}(t)$ for each content $i$, where $D_{i}(t)$ is the confidence interval quantifying the uncertainty in $p_{i}(t)$ at time $t$. We denote $\bar{w}^{\text{rec}}$ the mean of $w_{u}^{\text{rec }}$ over all user $u$. The confidence interval is defined as follows
$D_{i}(t)=\sqrt{\frac{2\left(1-\bar{w}^{\text{rec}}\right)^{1/\eta}\log t}{n_{i}(t)}+\frac{\alpha\left(\bar{w}^{\text{rec}}\right)}{2n_{i}(t)}}.$ (11)
The UCB-based algorithm chooses the top $C$ contents with the largest index and recommend any $R$ contents out of them.
Remark 1.
We have chosen such a confidence interval because we want the confidence interval to be low when $\bar{w}_{u}^{\text{rec}}\approx 1$. This is because users always tend to accept the recommendation when $\bar{w}_{u}^{\text{rec}}\approx 1$. Since we know distribution over recommended contents, which is the uniform distribution, there is less uncertainty about $p_{i}$’s. Similarly, the users ignore the recommendation when $\bar{w}_{u}^{\text{rec}}\approx 0$. In this case, the users follow their preference distributions $p_{u}^{\text{pref}}(i)$ ignoring the recommendations. Then, Problem (4) becomes standard CMAB without recommendation. Our algorithm is formally presented as Algorithm 1 below.
Algorithm 1 UCB Experiment with Recommendation
0:  $C$: Cache Capacity, $R$: Recommended Contents,Set of contents $\mathcal{N}=\{1,2,3,...,N\}$
0:  Updated statistics $n_{i}(t)$ and $p_{i}(t)$
1:  Initialization: $n_{i}(0)=1,p_{i}(0)=0\quad\forall i\in\mathcal{N}$
2:  for $t=1,2,\ldots,T$ do
3:     Compute UCBs
$U_{k}(t)=p_{k}(t)+D_{k}(t)$
for all $k\in\mathcal{N}$ using (10) and (11).
4:     Pick Top $C$ contents with the highest UCBs, $C(t)$, to cache.
5:     For each user, pick any $R$ contents from the cached contents to recommend
6:     Observe the number of requests for cached contents
7:     Update statistics: for all $i\in C(t)$,
$n_{i}(t)=n_{i}(t-1)+1$
$p_{i}(t)=\frac{p_{i}(t-1)n_{i}(t-1)+\text{no. of req for }i\text{ at time }t}{n_{i}(t)}$
8:  end for
Remark 2.
We have assumed the recommendation induced preference distributions $p_{u}^{rec}$’s to be uniform distributions (see (2)). However, the proposed UCB-based algorithm and the regret bound also apply to more general induced preference distributions wherein preference for a content depends depends only on the number of recommended contents $R$ and the tagged content’s position in the recommendation list. The key observation is that the optimal set of cached contents continues to be characterized by (8). Several works have modeled the induced preference distributions as Zipf distribution, given by (see [10, 17]).
$p_{u}^{\text{rec }}(i)=\frac{x_{ui}\sum_{k=1}^{R}l_{uik}k^{-\beta_{u}}}{\sum_{j=1}^{R}j^{-\beta_{u}}}$ (12)
where $\beta_{u}$ is the Zipf distribution parameter associated with user $u$ and $l_{uik}\in\{0,1\}$ are defined as follows.
$l_{uik}=\begin{cases}1&\text{if content $i$ is at the $k^{\text{th}}$ positionin user $u$'s}\\ &\text{recommendation list,}\\ 0&\text{otherwise.}\end{cases}$
For $\beta_{u}=0$ Zipf distribution reduces to the uniform distribution in (2). In Section IV, we evaluate the cache hit performance of the proposed algorithm for both uniform and Zipf distributions.
Remark 3.
Our proposed algorithm suggests that we recommend only the cached contents. In general, recommendations to users should also depend on their preferences. In particular, they are expected not to be far from the users’ preferences. This can be captured through following constraints on recommendations (see [5]).
$\sum_{i\in\mathcal{N}}x_{ui}p_{u}^{pref}(i)\geq Q_{u},\quad\forall u\in\mathcal{U},$ (13)
where $Q_{u}$ are user-specific thresholds. The quantities on the left in (13) are referred to as recommendation qualities. Incorporating these constraints in recommendation decisions is a part of our future work.
The following Lemma (1) is used in the proof of Theorem (1). This is another equivalent regret definition for (9).Let us define $\Delta_{e,k}\coloneqq\hat{p_{e}}-\hat{p_{k}}\quad\forall e\in\mathcal{Y},k\in\mathcal{N}\setminus\mathcal{Y}$.
Lemma 1.
The regret $\operatorname{R}(T)$ (9) can also be written as follows
$\operatorname{R}(T)=T\sum_{i\in\mathcal{Y}}\hat{p_{i}}-\sum_{t=1}^{T}\sum_{i\in y(t)}\hat{p_{i}}$
where $\hat{p_{i}}=\sum_{u}\left(1-w_{u}^{\text{rec}}\right)p_{u}^{\text{pref }}(i)$.
Proof.
Please see Appendix -A
The following theorem characterizes the regret of Algorithm 1.
Theorem 1.
The expected regret of Algorithm 1 is bounded as
$E[\operatorname{R}(T)]$ $\leqslant\sum_{k\in\mathcal{N}\setminus\mathcal{Y}}\left[\frac{16\left(1-\bar{w}^{\text{rec}}\right)^{1/\eta}\log T}{\Delta_{\min,k}}+\frac{4\alpha\left(\bar{w}^{\text{rec}}\right)}{\Delta_{\min,k}}\right]$
$\quad+\sum_{k\in\mathcal{N}\setminus\mathcal{Y}}\sum_{e\in\mathcal{Y}}\Delta_{e,k}\left(\frac{2e^{-\alpha\left(\bar{w}^{\text{rec}}\right)}}{4\left(1-\bar{w}^{\text{rec}}\right)^{1/\eta}-1}\right)$
for $\bar{w}^{\text{rec}}\leqslant 1-\frac{1}{4^{\eta}}$, where $\Delta_{\min,k}=\min_{e\in\mathcal{Y}}\Delta_{e,k}$, $\alpha>0$, and $\eta>0$.
Remark 4.
If we use Algorithm 1 without recommendation, i.e., setting $\bar{w}_{u}^{\text{rec}}=0$ for all $u$, then we get the standard CMAB. An upper bound on its regret can be derived following similar approach as in [9, Theorem 2]. It turns out to be as follows.
$\operatorname{R}(T)\leq\sum_{e\in\mathcal{N}\setminus\mathcal{Y}}\frac{16}{\Delta_{\min,e}}\log T+\sum_{k\in\mathcal{N}\setminus\mathcal{Y}}\sum_{e\in\mathcal{Y}}\Delta_{e,k}\frac{4}{3}\pi^{2}$
Theorem  1 offers a lower regret bound. Expectedly, recommending popular contents while caching these contents helps in achieving lower regret.
Remark 5.
The proof of Theorem 1 is motivated by [9, Theorem 2]. However, we use different confidence intervals, which warrant different arguments at several steps.
Proof.
We define the gaps $\Delta_{e,k}=\hat{p_{e}}-\hat{p_{k}}$ where $e\in\mathcal{Y},k\in\mathcal{N}\setminus\mathcal{Y}$ We also consider bijections $\pi^{y(t)}:y(t)\mapsto\mathcal{Y}$ satisfying $\pi^{y(t)}(k)=k$ for all $k\in y(t)\cap\mathcal{Y}$, which we use for an equivalent expression of the regret. We can express these bijections in terms of the following indicator functions.
$\mathds{1}_{e,k}(t)=\mathds{1}\left\{k\in y(t),\pi^{y(t)}(k)=e\right\}.$ (14)
From Lemma (1) , we have the following
$\operatorname{R}(T)$ $=T\sum_{i\in Y}\hat{p_{i}}-\sum_{t=1}^{T}\sum_{i\in y(t)}\hat{p_{i}}$
$=\sum_{t=1}^{T}\left[\sum_{i\in\mathcal{Y}}\hat{p_{i}}-\sum_{i\in y(t)}\hat{p_{i}}\right]$
$=\sum_{t=1}^{T}\sum_{k\in\mathcal{N}\setminus\mathcal{Y}}\sum_{e\in\mathcal{Y}}\Delta_{e,k}\mathds{1}_{e,k}(t)$ (15)
where the last equality is obtained using the definitions of the indicator functions as in (14). Taking expectation on both sides of (15), we get
$\mathbb{E}[\operatorname{R}(T)]=\sum_{k\in\mathcal{N}\setminus\mathcal{Y}}\sum_{e\in\mathcal{Y}}\Delta_{e,k}\mathbb{E}\left[\sum_{t=1}^{T}\mathds{1}_{e,k}(t)\right].$ (16)
Now we bound the expected cumulative regret associated with each content $k\in\mathcal{N}\setminus\mathcal{Y}$. The key idea of this step is to decompose the indicator function into two parts as follows
$\mathds{1}_{e,k}(t)=$ $\mathds{1}_{e,k}(t)\mathds{1}\left\{n_{k}(t-1)\leqslant l_{e,k}\right\}+$
$\mathds{1}_{e,k}(t)\mathds{1}\left\{n_{k}(t-1)>l_{e,k}\right\}$ (17)
where the numbers $l_{e,k}$ are chosen as explained below. We rewrite (16) using (17) as follows
$\mathbb{E}[\operatorname{R}(T)]=\sum_{k\in\mathcal{N}\setminus\mathcal{Y}}\sum_{e\in\mathcal{Y}}\Delta_{e,k}\mathbb{E}\left[\sum_{t=1}^{T}\mathds{1}_{e,k}(t)\mathds{1}\left\{n_{k}(t-1)\leqslant l_{e,k}\right\}\right]$
$\quad+\sum_{k\in\mathcal{N}\setminus\mathcal{Y}}\sum_{e\in\mathcal{Y}}\Delta_{e,k}\mathbb{E}\left[\sum_{t=1}^{T}\mathds{1}_{e,k}(t)\mathds{1}\left\{n_{k}(t-1)>l_{e,k}\right\}\right].$ (18)
Then we bound each term of (III) using Lemma (2) and (3).
Let us define sets $\tau_{i}\coloneqq\{t\leq T:i\text{ is cached and recommended at time }t\}$ for all $i$. We can then estimate $\hat{p_{i}}$’s using $\hat{p_{i}}(s)$’s defined as follows.
$\hat{p_{i}}(t)$ $\coloneqq\frac{\sum_{s\leq t}Z_{i}(s)-\sum_{s\in\tau_{i}}\sum_{u}\frac{w_{u}^{\text{rec}}}{R}}{n_{i}(t)}$
$=\frac{\sum_{s=1}^{t}\bigg{[}Z_{i}(s)-\sum_{u}\frac{w_{u}^{\text{rec}}}{R}\mathds{1}\{s\in\tau_{i}\}\bigg{]}}{n_{i}(t)}.$
Further, recall that $\mathds{1}_{e,k}(t)=1$ implies that $k\in y(t)$ and $e=k$ or $e\in\mathcal{Y}\setminus y(t)$. In either case, $U_{k}(t)\geqslant U_{e}(t)$ which in turn implies at least one of the following three events.
  1. 1.
    $\hat{p_{e}}(t)\leqslant\hat{p_{e}}-D_{e}(t)\quad\forall e\in Y$
  2. 2.
    $\hat{p_{k}}(t)\geqslant\hat{p_{k}}+D_{k}(t)$
  3. 3.
    $\Delta_{e,k}\leqslant 2D_{k}(t).$
We can bound the probabilities of Events $1$ and $2$ using Hoeffoding’s inequalities as follows.
$P\bigg{[}\hat{p_{k}}(t)-\hat{p_{k}}$ $\geqslant D_{k}(t)\bigg{]}\leqslant\frac{e^{-\alpha\left(\bar{w}^{\text{rec}}\right)}}{t^{4\left(1-\bar{w}^{\text{rec}}\right)^{1/\eta}}}$ (19)
$P\bigg{[}\hat{p_{e}}(t)-\hat{p_{e}}$ $\geqslant D_{k}(t)\bigg{]}\leqslant\frac{e^{-\alpha\left(\bar{w}^{\text{rec}}\right)}}{t^{4\left(1-\bar{w}^{\text{rec}}\right)^{1/\eta}}}.$ (20)
If $n_{k}(t-1)>4\left(\frac{2\left(1-\bar{w}^{\text{rec}}\right)^{1/\eta}\log T}{\Delta_{e,k}^{2}}+\frac{\alpha\left(\bar{w}^{\text{rec}}\right)}{2\Delta_{e,k}^{2}}\right)$, rearranging the terms gives
$\Delta_{e,k}>2D_{k}(t),$
and so, Event 3 cannot happen in this case. Hence we choose
$l_{e,k}=\left\lceil 4\left(\frac{2\left(1-\bar{w}^{\text{rec}}\right)^{1/\eta}\ \log T}{\Delta_{e,k}^{2}}+\frac{\alpha\left(\bar{w}^{\text{rec}}\right)}{2\Delta_{e,k}^{2}}\right)\right\rceil.$ (21)
Now, using (19) and (20), and choosing $l_{e,k}$ as stated above, we can state the following Lemmas as follows.
Lemma 2.
$\sum_{e\in\mathcal{Y}}$ $\Delta_{e,k}\mathbb{E}\bigg{[}\sum_{t=1}^{T}\mathds{1}_{e,k}(t)\mathds{1}\left\{n_{k}(t-1)>l_{e,k}\right\}\bigg{]}$
$\leq\sum_{e\in\mathcal{Y}}\Delta_{e,k}\left(\frac{2e^{-\alpha\left(\bar{w}^{\text{rec}}\right)}}{4\left(1-\bar{w}^{\text{rec}}\right)^{1/\eta}-1}\right)$ (22)
Proof.
Please see Appendix -C
Lemma 3.
$\sum_{e\in\mathcal{Y}}\Delta_{e,k}\mathbb{E}\bigg{[}\sum_{t=1}^{T}\mathds{1}_{e,k}(t)\mathds{1}\left\{n_{k}(t-1)\leqslant l_{e,k}\right\}\bigg{]}$
${\leqslant}\frac{16\left(1-\bar{w}^{\text{rec}}\right)^{1/\eta}\log T}{\Delta_{\min,k}}+\frac{4\alpha\left(\bar{w}^{\text{rec}}\right)}{\Delta_{\min,k}}$ (23)
Proof.
Please see Appendix -B
Now using Lemma (2), (3) in (III) and further summing over $k\in\mathcal{N}\setminus\mathcal{Y}$, we get the regret bound in Theorem 1. ∎
Figure 1: Performace of our algorithm
Figure 2: Performace of our algorithm for various $w_{u}^{rec}$
Figure 3: Performace of our algorithm for various $\alpha$
Figure 4: Performace of our algorithm
IV Numerical Results
In this section, we numerically evaluate the UCB-based algorithm (1) for various system parameters. We will compare the performance of our algorithm with the standard CMAB-UCB algorithm [9], the greedy algorithm, $\epsilon$-Greedy algorithm. In all the experiments, we assume a total number of contents $N=50$, cache capacity $C=20$ and the number of users $U=20,\eta=4$.
Performance of Algorithm (1) with CMAB-UCB [9], $\epsilon$-Greedy and Greedy algorithm with uniform distribution among the recommended contents
We compare the performance of Algorithm (1) with CMAB-UCB ,$\epsilon$-Greedy and Greedy algorithm. In the Greedy algorithm, we calculate $p_{i}(t)$ for all $i\in\mathcal{N}$. Then, pick the top $C$ contents with the highest $p_{i}(t)$. In the $\epsilon$-Greedy algorithm, a random number is generated every time. If it exceeds the epsilon, the top $C$ contents with the highest $p_{i}(t)$ are selected. Else $C$ contents selected uniformly from content set $\mathcal{N}$. In this experiment, we used $w_{u}^{rec}=0.95$ for all $u\in\mathcal{U}$, $\alpha=5,\epsilon=0.4$. We can see from the figure Fig.(1) that our algorithm performs better compared to better to other algorithms.
Performance of Algorithm (1) with CMAB-UCB [9] for various $w_{u}^{rec}$
In this experiment will compare the performance of our algorithm for various $w_{u}^{rec}$ values i.e. $w_{u}^{rec}=0.99,0.95,0.9$ for all $u\in\mathcal{U}$. When $w_{u}^{rec}\approx 1$, users always tend to accept the recommendation. Since we recommending cache contents, we expect the regret to be low. Hence, we want regret to decrease when $w_{u}^{rec}$ increases and becomes close to 1. This can be seen from the Fig.(2)
Performance of Algorithm (1) with CMAB-UCB [9] for various $\alpha$
In this experiment we will compare the performance of our algorithm for various $\alpha$ values i.e. $\alpha=1,5,10$ keeping $w_{u}^{rec}=0.95$ for all $u\in\mathcal{U}$. From the Fig.(3), we can see that regret does not vary much with $\alpha$.
Performance of Algorithm (1) with CMAB-UCB [9], $\epsilon$-Greedy and Greedy algorithm with Zipf distribution among the recommended contents
In this experiment , we assume Zipf distribution (12) over the recommended contents instead of uniform distribution. This makes the probability of requests to depend on the position of content in the recommended list. Here we also assume $w_{u}^{\text{rec}}$ is different for $u\in\mathcal{U}$ and are choosen uniformly from $[0.9,0.99]$ for each user. Similarly $\beta_{u}$ is choosen uniformly from $[1,2]$ for each user. Also $\alpha$ is fixed at $5$. Then we compare the performance of our algorithm for both uniform and Zipf distribution over recommended contents with the other algorithms and is shown in the Fig.(4). From the Fig.(4), we can see that having Zipf distribution over the recommended contents yields the same performance as uniform distribution as we justified in Remark(2).
V Conclusion and Future Work
In this work, we analysed the content caching problem with the recommendations, modelling it as a CMAB. We proposed a UCB-based algorithm and also provided an upper bound on the regret of the algorithm. We then compared the performance of this algorithm to that of a few existing algorithms. We also considered the case where the position of the recommended content affects its request probability. We found that caching with recommendation improves the cache hit performance in all the cases.
In our future work, we would like to consider a different recommendation model where each user examines the contents in the recommendation list one by one until he finds attractive content, and once he finds any attractive content, he does not examine the rest. We would like to incorporate recommendation quality (13) into our problem and would also like to consider the case where $w_{u}^{\mathrm{rec}}$’s are unknown.